OpenStudy (anonymous):

WILL GIVE A MEDAL What is the equation of a line that passes through point (1, -2) and has a slope of 1/3?

jimthompson5910 (jim_thompson5910):

y = mx+b is the general equation m is the slope b is the y-intercept

jimthompson5910 (jim_thompson5910):

m = 1/3 is given

jimthompson5910 (jim_thompson5910):

(1,-2) is the given point so x = 1 and y = -2

jimthompson5910 (jim_thompson5910):

y = mx+b y = (1/3)*x+b -2 = (1/3)*(1)+b solve for b

OpenStudy (anonymous):

But I got -7/3, but the answer choices are 3x - y - 7 = 0 x - 3y + 7 = 0 x - 3y - 7 = 0

OpenStudy (anonymous):

if use the more general form y=y1=m(x-x1) that would easier y-(-2)=1/3(x-1)

OpenStudy (anonymous):

*(y-y1)=m(x-x1)

OpenStudy (anonymous):

and then solve for y

OpenStudy (anonymous):

okay just multbly the whole equation by 3

OpenStudy (anonymous):

i hope that is clear now for you @z053sme